# Controllable Emotion Generation with Emotion Vectors

### v1

@@ C.1 Perplexity
$\text{Perplexity}=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P(y_{i}|y_{1:i-1})\right)$
@@ C.2 Topic adherence
TA=Number of adherent sentencesTotal number of $\text{TA}=\frac{\text{Number of {adherent} sentences}}{\text{Total number of sentences}}$ (2)
C.3 Emotion Probability Score
@@
The bart-large-mnli model is specifically designed for natural language understanding tasks, particularly natural language inference and zero-shot text classification. By leveraging the extensive pre-training of BART combined with the diverse and comprehensive MNLI dataset, facebook/bart-large-mnli is capable of effectively determining the relationship between sentence pairs, such as entailment, contradiction, and neutrality. Its robust performance in zero-shot classification tasks makes it a valuable tool for applications requiring flexible and accurate text classification without the need for task-specific training data. Additionally, the model’s ability to handle custom labels allows us to tailor the classification process to our specific needs, ensuring that the emotional intensity of generated text is accurately and effectively measured. To evaluate the emotional intensity of the generated sentences, we input each sentence produced by our models into the facebook/bart-large-mnli classifier. For example, consider the sentence: "I get so scared! My heart races, I can’t breathe, and I just want to hide." This sentence is directly fed into the model, which then classifies it into one of the three predefined categories: emotionless, neutral, or emotional.
@@
EPR=Number of emotional classificationsTotal number of $\text{EPR}=\frac{\text{Number of {emotional} classifications}}{\text{Total number of sentences}}$ (3)
@@
C.4 Emotion Absolute Score
@@
C.5 Visualization of Emotion Vectors

### v2

@@ D.1 Perplexity
$\text{Perplexity}=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P(y_{i}|y_{1:i-1})\right)$ (55)
@@ D.2 Topic adherence
$\text{TA}=\frac{\text{Number of {adherent} sentences}}{\text{Total number of sentences}}$ (56)
D.3 Emotion Probability Score
@@
The bart-large-mnli model is specifically designed for natural language understanding tasks, particularly natural language inference and zero-shot text classification. By leveraging the extensive pre-training of BART combined with the diverse and comprehensive MNLI dataset, facebook/bart-large-mnli is capable of effectively determining the relationship between sentence pairs, such as entailment, contradiction, and neutrality. Its robust performance in zero-shot classification tasks makes it a valuable tool for applications requiring flexible and accurate text classification without the need for task-specific training data. Additionally, the model’s ability to handle custom labels allows us to tailor the classification process to our specific needs, ensuring that the emotional intensity of generated text is accurately and effectively measured. To evaluate the emotional intensity of the generated sentences, we input each sentence produced by our models into the facebook/bart-large-mnli classifier. For example, consider the sentence: ”I get so scared! My heart races, I can’t breathe, and I just want to hide.” This sentence is directly fed into the model, which then classifies it into one of the three predefined categories: emotionless, neutral, or emotional.
@@
$\text{EPR}=\frac{\text{Number of {emotional} classifications}}{\text{Total number of sentences}}$ (57)
@@
D.4 Emotion Absolute Score
@@
D.5 Target Emotion Confidence
D.5.1 Computation of Target Emotion Confidence (TEC)
To quantitatively evaluate how well the generated response aligns with the desired target emotion, we introduce the Target Emotion Confidence (TEC) score. This score reflects the degree of emotional alignment based on external classification.
Classifier Details
We adopt the facebook/bart-large-mnli model as an external emotion classifier. This model is a BART-based transformer fine-tuned on the Multi-Genre Natural Language Inference (MNLI) dataset. It is widely used for zero-shot or prompt-based classification tasks due to its robust generalization. In our setup, we adapt the classifier to perform emotion recognition over six emotion classes: anger, disgust, fear, joy, sadness, and neutral.
Multi-label Classification
Unlike standard single-label classification, we use a multi-label formulation where each generated response is assigned a probability for every emotion label independently. This setting reflects the fact that emotional content can have overlapping characteristics and avoids forcing an exclusive prediction.
TEC Score Definition
Let $\mathcal{R}_{m,e}^{(\lambda)}$ be the set of responses generated by model $m$ when applying EV of emotion $e$ at intensity $\lambda\in\{1,2,4\}$ on the EQ+ dataset. Let $C(r,e)$ be the classifier’s predicted probability for target emotion $e$ given response $r$. Then, the TEC score is defined as:
$\mathrm{TEC}(m,e,\lambda)=\frac{1}{|\mathcal{R}_{m,e}^{(\lambda)}|}\sum_{r\in\mathcal{R}_{m,e}^{(\lambda)}}C(r,e)$ (59)
This score reflects the average classifier confidence that the generated responses express the intended target emotion.
Example
For instance, to compute the TEC score for model LLaMA2-7B under 2× anger EV, we:
  • Apply the 2× anger EV to LLaMA2-7B across all EQ+ prompts;
  • Collect the generated responses;
  • Pass each response through the classifier and extract the probability for anger;
  • Average these probabilities.
This process is repeated across models, emotions, and EV intensities. The resulting scores has been reported in Table 4.
D.5.2 TEC Matrices for Emotionally Biased Prompts
Table 6 presents six TEC score matrices, each corresponding to a distinct target emotion. These scores are computed on the emotionally biased subset of the EQ+ dataset using the Qwen-2.5 model, as described in Section 4.X.
For each target emotion, we evaluate the impact of applying EVs at different intensities (0×, 1×, 2×, 4×) on prompts originally designed to express a specific emotion (rows). The values in each matrix represent the average Target Emotion Confidence (TEC) score for the specified EV setting.
These results demonstrate that even when queries are emotionally suggestive, the EV mechanism is able to effectively shift the emotional output of the model. Stronger EV intensities generally produce higher TEC scores, confirming the controllability of emotional expression via EVs.
Target Emotion: Anger Target Emotion: Disgust
[…]
anger 37.09 74.68 97.18 98.43 anger 18.74 44.73 93.76 94.42
disgust 16.95 68.30 97.35 93.70 disgust 25.48 81.04 94.69 91.87
fear 15.66 35.84 95.38 94.67 fear 11.24 16.42 93.76 96.59
joy 0.34 1.15 92.21 96.09 joy 0.15 0.08 81.58 91.98
sadness 10.36 44.77 92.21 96.35 sadness 6.28 12.94 85.19 93.04
neutral 10.56 14.06 94.93 95.40 neutral 7.30 9.99 92.31 91.18
Target Emotion: Fear Target Emotion: Joy
[…]
anger 33.21 63.59 94.89 95.56 anger 24.81 73.34 96.37 67.58
disgust 19.79 59.77 93.84 94.14 disgust 9.79 64.85 96.30 71.92
fear 50.83 86.60 94.95 91.96 fear 17.30 68.93 92.39 63.64
joy 0.98 6.61 80.08 95.37 joy 66.29 90.52 94.61 63.01
sadness 14.25 62.16 93.88 97.13 sadness 14.31 59.00 94.30 62.54
neutral 12.55 16.29 83.42 90.60 neutral 25.77 46.33 90.59 52.71
[…]
Table 6: TEC scores under different EV intensities for each target emotion. Each subtable corresponds to a specific target emotion, indicating the type of Emotion Vector (EV) applied during generation. Rows represent the original emotion label of the query in the EQ+ dataset, and columns denote the EV intensity (i.e., 0×, 1×, 2×, 4×). The values in each cell reflect the classifier-assigned probability that the generated response expresses the target emotion. This structure allows us to examine how increasing the strength of a specific EV influences the emotional expression of the model, even when the input query is emotionally biased toward a different category. As shown, applying stronger EVs leads to substantial gains in target emotion alignment for non-matching queries, demonstrating the controllability and robustness of our EV-based generation framework.